A train leaves Kansas City for Dallas and it travels at an average speed of 60 miles/hour. Another train leaves in the opposite direction and travels at an average speed of 80 miles/hour. At what point [with respect to either of the stations] would they meet?
Answer by ankor@dixie-net.com(22740)

A train leaves Kansas City for Dallas and it travels at an average speed of 60 miles/hour.
Another train leaves in the opposite direction and travels at an average speed of 80 miles/hour.
At what point [with respect to either of the stations] would they meet?
:
Assuming 454 mi from Dallas to KC
Let t = time for them to meet, when they meet the total distance will be 454 mi
:
60t + 80t = 454
140t = 454
t = 454/140
t = 3.243 hrs
:
Find the distance from KC
3.243 * 60 = 194.6 mi they will meet